OpenStudy (anonymous):

Show that for every integer n>1, there are integers x > y>1 such that \[ \frac { x^2 + y^2 -1}{ x y} =n \]

OpenStudy (anonymous):

Are there more conditions on x and y? Because you could always take x = n and y = 1. That would work for any integer n.

OpenStudy (anonymous):

bah i take that back. y must be greater than 1.

OpenStudy (kinggeorge):

I'm back for a moment to propose that using \(x=n\) and \(y=n^2-1\) has been working for the examples I've looked at so far. Now I'm actually off to bed :P

OpenStudy (anonymous):

yeah, thats the solution:\[\frac{x^2+y^2-1}{xy}=\frac{n^2+(n^2-1)^2-1}{n(n^2-1)}=\frac{n^2+n^4-2n^2+1-1}{n(n^2-1)}\]\[=\frac{n^4-n^2}{n(n^2-1)}=\frac{n^2(n^2-1)}{n(n^2-1)}=n.\] Nice insight :)

OpenStudy (anonymous):

Maybe just switch x and y so that the x > y > 1 statement holds.